2. Calculate the pH to the 100th's place when 10mL of 0.890M ammonium iodide NH4I (pka=9.26) is titrated with 1.436 M sodium hydroxide to the equivalence point.

Can an expert help me solve the question above?
Answer #1

Given: pka of NHAI : 9.26 Concentration of NHI 0.8901 Volume of NHA I = lomb =0.012 Concentration of NaOH = 1.436 M SaleebianNeember of indes og NHL I = Moleky x volume =(0.890 melly) (0.014) Muit = 0.0089 mil RHLE Z Based on a balanced chemical equaNha ay) +4%) = Memo) + Pers) Weak base Concentaation - males of NH₃ Total volume 13 Total volume call +0.0062 h = 0.01622 0.0Dissociation constant of bass as the ratio of the product of the products to the product of the reactants each paired to coof| POH= 2.50 pH + pOH = 14 pH = 14 - POH pH = 14-2.50 [ pH = 11.50 Ones tenths tens & 11.50 Hundredths decimal point d the Bol

pH of a solution is 11.50